# Mamba4Rec: Towards Efficient Sequential Recommendation with Selective State Space Models

### v1

@@ Abstract.
Sequential recommendation aims to estimate the dynamic user preferences and sequential dependencies among historical user behaviors. Although Transformer-based models have proven to be effective for sequential recommendation, they suffer from the inference inefficiency problem stemming from the quadratic computational complexity of attention operators, especially for long-range behavior sequences. Inspired by the recent success of state space models (SSMs), we propose Mamba4Rec, which is the first work to explore the potential of selective SSMs for efficient sequential recommendation. Built upon the basic Mamba block which is a selective SSM with an efficient hardware-aware parallel algorithm, we incorporate a series of sequential modeling techniques to further promote the model performance and meanwhile maintain the inference efficiency. Experiments on two public datasets demonstrate that Mamba4Rec is able to well address the effectiveness-efficiency dilemma, and defeat both RNN- and attention-based baselines in terms of both effectiveness and efficiency. The code is available at https://github.com/chengkai-liu/Mamba4Rec.
@@ 1. Introduction
Personalized online services rely heavily on sequential recommender systems to capture dynamic user preferences (wang2021survey; lu2015recommender; lin2023map). These systems aim to predict users’ future interactions by effectively modeling sequential dependencies among historical user behaviors. Earlier neural approaches adopt Convolutional Neural Networks (CNNs) (tang2018personalized) and Recurrent Neural Networks (RNNs) (hidasi2015session; li2017neural) which have pioneered the application of neural networks in sequential recommendation yet suffer from the catastrophic forgetting issue (schak2019study; kirkpatrick2017overcoming). More recently, researchers start to introduce Transformer (vaswani2017attention) as the backbone module and leverage the self-attention mechanism to model the dynamic user behavior sequences (kang2018self; yuan2022multi). In spite of their remarkable performance, these attention-based methods generally suffer from the inference inefficiency problem due to the quadratic computational complexity inherent in attention operators, especially for long user behavior sequences or large-scale datasets.
@@
To address the dilemma of recommendation performance and inference efficiency, in this paper, we consider the state space model (SSM) (gu2021efficiently) as the core operator for sequential recommendation. SSMs (e.g., S4 (gu2021efficiently), S5 (smith2022simplified), Mega (ma2022mega)) have been widely adopted as alternatives to RNNs, CNNs, and Transformers in various natural language tasks. They inherently possess inference efficiency due to their recurrent nature, and meanwhile enjoy strong performance to handle long-range dependencies through structured state matrices. Moreover, a recently proposed SSM variant (i.e., Mamba (gu2023mamba)) further introduces the data-dependent selective mechanism with efficient hardware-aware design to tackle the data- and time-invariant issue of prior SSMs. This allows the model to selectively extract essential knowledge and filter out irrelevant noise according to the input data, thus leading to superior sequential modeling performance.
@@
In this work, we propose Mamba4Rec, which is the first work to explore the potential of selective SSM for efficient sequential recommendation. Built upon the basic Mamba block (gu2023mamba), we incorporate and discuss a series of techniques (e.g., position embedding (vaswani2017attention), residual connection (he2016deep), layer normalization (ba2016layer), and position-wise feed-forward network (vaswani2017attention)), which supplements the system non-linearity, stabilizes the training dynamic, and thereby further promotes the sequential modeling capability.
The main contributions of this paper are as follows:
@@
  • We are the first to explore the potential of selective SSMs for sequential recommendation, which addresses the dilemma of recommendation performance and inference efficiency.
@@
  • We propose Mamba4Rec by incorporating the basic Mamba block with various techniques like layer normalization and feed-forward networks, which further improves the sequential modeling capability without sacrificing the inference efficiency.
@@
  • Experiments on two public datasets demonstrate the superiority of our proposed Mamba4Rec compared with RNN- and attention-based baselines in terms of both effectiveness and efficiency.
@@ 2.1. Problem Statement
2.2. State Space Models
The State Space Model (SSM) is a sequence modeling framework based on linear ordinary differential equations. It maps an input sequence $x(t)\in\mathbb{R}^{D}$ to the output sequence $y(t)\in\mathbb{R}^{N}$ through the latent state $h(t)\in\mathbb{R}^{N}$:
@@ 2.2. State Space Models
where $\overline{\bm{A}}=\exp(\Delta\bm{A})$ and $\overline{\bm{B}}=(\Delta\bm{A})^{-1}(\exp(\Delta\bm{A})-\bm{I})\cdot\Delta\bm{B}$. After converting the parameters from a continuous form $(\Delta,\bm{A},\bm{B},\bm{C})$ to a discrete form $(\overline{\bm{A}},\overline{\bm{B}},\bm{C})$, the model can be computed in a linear recurrence way (gu2021combining), which enhances computational efficiency. Deriving from the vanilla SSM, the structured state space model (S4) (gu2021efficiently) imposes structure on the state matrix $\bm{A}$ with HiPPO (gu2020hippo) initialization to further improve the long-range dependency modeling.
Building upon S4, Mamba (gu2023mamba) introduces a data-dependent selection mechanism and leverages a hardware-aware parallel algorithm in recurrent mode. This combined approach empowers Mamba to capture contextual information effectively, particularly in long sequences, while maintaining computational efficiency. As a linear-time sequence model, Mamba achieves Transformer-quality performance with better efficiency, especially on long sequences.
@@ 3.2. Embedding Layer
Similar to existing models, our approach utilizes an embedding layer to map item IDs to a high-dimensional space. This embedding layer uses a learnable embedding matrix $\bm{E}\in\mathbb{R}^{|\mathcal{V}|\times D}$, where $D$ is the embedding dimension. By applying the embedding layer to the input item sequence $\mathcal{S}_{u}$, we obtain the item embeddings $H$. To enhance robustness and prevent overfitting, we incorporate both embedding dropout and layer normalization after retrieving the embeddings:
(3) $H$ $=\operatorname{LayerNorm}\left(\operatorname{Dropout}\left(H_{o}\right)\right)\in\mathbb{R}^{n_{u}\times D}.$
During the training and inference stage, we merge multiple samples into one mini-batch, resulting the input data $x\in\mathbb{R}^{B\times L\times D}$ for the subsequent Mamba block, where $B$ is the batch size and $L$ is padded sequence length.
@@ 3.3. Mamba Layer
Table 1. Recommendation performance. The best results are bold, and the second-best are underlined.
@@
Mamba Block. As detailed in Algorithm 1, the Mamba block operates on an input $x_{t}\in\mathbb{R}^{B\times L\times D}$ with dimensions of batch size $B$, sequence length $L$, and hidden dimension $D$ at each time step $t$. It begins by performing linear projections on the input $x_{t}$ with an expanded hidden dimension of $D$ to obtain $x$ and $z$. These projections are then processed through a 1D convolution and a SiLU (elfwing2018sigmoid) activation. The core of the block involves a selective state space model with parameters discretized based on the input. This discretized SSM, along with $x^{\prime}$, generates the state representation $y$. Finally, $y$ is combined with a residual connection from $z$ after applying SiLU, and a final linear projection delivers the final output $y_{t}$ at time step $t$. Overall, the Mamba block leverages input-dependent adaptations and selective SSM to process sequential information effectively. The parameters of the Mamba block consist of an SSM state expansion factor $N$, a kernel size $K$ for the convolution, and a block expansion factor $E$ for input and output linear projections. Selecting the expansion factors $N$ and $E$ involves a tradeoff between (capturing complex relationships) and training efficiency (higher $N,E$ increase computational cost).
The Mamba block addresses the prevalent effectiveness-efficiency dilemma existing in RNNs and Transformers. The recurrent SSM in the Mamba block operates in linear time with respect to the sequence length, making it as efficient as RNNs for inference, achieved by discretized parameters. Additionally, the selection mechanism makes the parameters of SSM data-dependent. This enables Mamba to selectively remember or forget information, crucial for capturing complex dependencies within sequential data. With the structured state matrix addressing long-range dependencies with HiPPO, it outperforms self-attention blocks in modeling long sequences.
@@
1:Input: $x_{t}:(B,L,D)$
2:Output: $y_{t}:(B,L,D)$
3:$x,z:(B,L,ED)\leftarrow\operatorname{Linear}(x_{t})$triangleright Linear projections of input
[…]
@@
7:$\Delta:(B,L,D)\leftarrow\operatorname{Softplus}(\text{Parameter}+\operatorname{Broadcast}(\operatorname{Linear}(x^{\prime})))$
@@
9:$y:(B,L,ED)\leftarrow\text{SelectiveSSM}(\overline{\bm{A}},\overline{\bm{B}},\bm{C})(x^{\prime})$
10:$y^{\prime}:(B,L,ED)\leftarrow y\otimes\operatorname{SiLU}(z)$
11:$y_{t}:(B,L,D)\leftarrow\operatorname{Linear}(y^{\prime})$triangleright Final linear projection
12:return $y_{t}$
@@ 3.4. Prediction Layer
(5) $\hat{y}=\operatorname{Softmax}\left(\operatorname{Linear}\left(h\right)\right)\in\mathbb{R}^{|\mathcal{V}|},$
@@ 4.1. Experimental Setup
Video-Games $14,494$ $6,950$ $132,209$ $9.1$
Baselines. We compare Mamba4Rec against several baseline methods, which include RNN-based models like GRU4Rec (hidasi2015session) and NARM (li2017neural), as well as Transformer-based models such as SASRec (kang2018self) and BERT4Rec (sun2019bert4rec).
@@
Implementation Details. In the default architecture of our model, a single Mamba layer is employed without position embeddings. The Adam optimizer (kingma2014adam) is used with a learning rate 0.001. The training batch size is 2048, and the evaluation batch size is 4096. All models use an embedding dimension of 64. To address the sparsity of the Amazon datasets, a dropout rate of 0.4 is used, compared to 0.2 for MovieLens-1M. The maximum sequence length is set proportionally to the mean number of actions per user: 200 for MovieLens-1M and 50 for Amazon datasets. We follow RecBole (zhao2021recbole) for further implementation details. For the parameters of the Mamba block, the SSM state expansion factor is 32, the kernel size for 1D convolution is 4, and the block expansion factor for linear projections is 2.
@@ 4.2. Overall Performance
Table 1 demonstrates the overall recommendation performance. Among sequential recommendation methods, Transformer-based models SASRec and BERT4Rec generally outperform RNN-based baselines. Notably, the state-space-based Mamba4Rec achieves superior performance compared to both RNN and Transformer-based baselines across sparse and dense datasets, as well as varying maximum sequence lengths. Compared to SASRec, Mamba4Rec achieves significantly better performance on the MovieLens-1M dataset, with longer sequence lengths. This highlights the potential of state space models as a compelling alternative to Transformers and RNNs architectures for sequential recommendation.
@@ 4.3. Model Efficiency
Table 3 details the efficiency performance of Transformer-based models and Mamba4Rec, all measured on a single Nvidia A5000 GPU. Considering the tradeoff between efficiency and effectiveness, we use the model with a single Mamba layer instead of models with only a Mamba block or stacked Mamba layers on the long-sequence MovieLens-1M dataset. Mamba4Rec exhibits superior performance in terms of GPU memory usage, training time per epoch, and inference time per batch. Additionally, it demonstrates faster convergence, reaching its best validation performance around epoch 50, compared to SASRec (epoch 70) and BERT4Rec (epoch 180). These results suggest that Mamba4Rec achieves superior efficiency and reduces GPU memory cost, particularly on long-sequence datasets when compared to Transformer-based models.
@@
Mamba4Rec 4.82GB 75.52s 0.14s
4.4. Ablation Study
Table 4. Ablation analysis on ML-1M and Beauty.

### v2

@@ Abstract.
Sequential recommendation aims to estimate the dynamic user preferences and sequential dependencies among historical user behaviors. Although Transformer-based models have proven to be effective for sequential recommendation, they suffer from the inference inefficiency problem stemming from the quadratic computational complexity of attention operators, especially for long behavior sequences. Inspired by the recent success of state space models (SSMs), we propose Mamba4Rec, which is the first work to explore the potential of selective SSMs for efficient sequential recommendation. Built upon the basic Mamba block which is a selective SSM with an efficient hardware-aware parallel algorithm, we design a series of sequential modeling techniques to further promote model performance while maintaining inference efficiency. Through experiments on public datasets, we demonstrate how Mamba4Rec effectively tackles the effectiveness-efficiency dilemma, outperforming both RNN- and attention-based baselines in terms of both effectiveness and efficiency. The code is available at https://github.com/chengkai-liu/Mamba4Rec.
@@ 1. Introduction
Personalized online services rely heavily on sequential recommender systems to capture dynamic user preferences (clickprompt; wang2021survey; lu2015recommender; lin2023map). These systems aim to predict users’ future interactions by effectively modeling sequential dependencies among historical user behaviors. Earlier neural approaches adopt convolutional neural networks (CNNs) (tang2018personalized) and recurrent neural networks (RNNs) (hidasi2015session; li2017neural), which have pioneered the application of neural networks in sequential recommendation yet suffer from the catastrophic forgetting issue (schak2019study; kirkpatrick2017overcoming). More recently, researchers have adopted the Transformer (vaswani2017attention) as the backbone module and leverage the self-attention mechanism to model the dynamic user behavior sequences (kang2018self; yuan2022multi). Despite their remarkable performance, these attention-based methods generally suffer from the inference inefficiency problem due to the quadratic computational complexity inherent in attention operators, especially for long user behavior sequences or large-scale datasets (tay2020long; tay2022efficient). This tradeoff between performance and efficiency has hindered their practical application.
@@
Recently, state space models (SSMs) (gu2021efficiently), such as S4 (gu2021efficiently), S5 (smith2022simplified) and Mega (ma2022mega), have been widely adopted as alternatives to RNNs, CNNs, and Transformers in various natural language tasks. These models inherently possess inference efficiency due to their recurrent nature, and meanwhile enjoy strong performance to handle long-range dependencies through structured state matrices. Moreover, a recently proposed SSM variant (i.e., Mamba (gu2023mamba)) further introduces the input-dependent selective mechanism with efficient hardware-aware design to tackle the data-independent issue of prior SSMs. This allows the model to selectively extract essential knowledge and filter out irrelevant noise according to the input data, thus leading to superior sequential modeling performance. These advancements provide the potential to address the tradeoff between recommendation performance and inference efficiency, positioning SSMs as a core operator for sequential recommendation.
@@
In this work, we provide an in-depth investigation of adopting selective SSMs for sequential recommendation, introducing Mamba4Rec, the first model to leverage the power of selective SSMs for efficient sequential recommendation. Built upon the basic Mamba block (gu2023mamba), we incorporate and discuss a series of techniques (e.g., position embedding (vaswani2017attention), residual connection (he2016deep), layer normalization (ba2016layer), and position-wise feed-forward network (vaswani2017attention)), which supplements the system non-linearity, stabilizes the training dynamics, and thereby further promotes the sequential modeling capability. The main contributions of this paper are as follows:
@@
  • We are the first to explore the potential of selective SSMs for sequential recommendation, to address the dilemma of recommendation performance and inference efficiency.
@@
  • By understanding the impact of the basic Mamba block, including its incorporation of normalization and feed-forward networks for sequential recommendation, we propose Mamba4Rec, which further improves sequential modeling capability without sacrificing inference efficiency.
@@
  • Experiments on public datasets demonstrate the superiority of our proposed Mamba4Rec compared with RNN- and attention-based baselines in terms of both effectiveness and efficiency.
@@ 2. Preliminaries
State Space Models (SSM). The State Space Model (SSM) is a sequence modeling framework based on linear ordinary differential equations. It maps an input sequence $x(t)\in\mathbb{R}^{D}$ to the output sequence $y(t)\in\mathbb{R}^{N}$ through the latent state $h(t)\in\mathbb{R}^{N}$:
@@
where $\overline{\bm{A}}=\exp(\Delta\bm{A})$ and $\overline{\bm{B}}=(\Delta\bm{A})^{-1}(\exp(\Delta\bm{A})-\bm{I})\cdot\Delta\bm{B}$. After converting the parameters from a continuous form $(\Delta,\bm{A},\bm{B},\bm{C})$ to a discrete form $(\overline{\bm{A}},\overline{\bm{B}},\bm{C})$, the model can be computed in a linear recurrence way (gu2021combining), which enhances computational efficiency. Deriving from the vanilla SSM, the structured state space model (S4) (gu2021efficiently) imposes structure on the state matrix $\bm{A}$ with HiPPO (gu2020hippo) initialization to further improve the long sequence modeling.
Building upon S4, Mamba (gu2023mamba) introduces an input-dependent selection mechanism and leverages a hardware-aware parallel algorithm in recurrent mode. This combined approach empowers Mamba to capture contextual information effectively, particularly in long sequences, while maintaining computational efficiency. As a linear-time sequence model, Mamba achieves Transformer-quality performance with better efficiency, especially on long sequences.
@@ 3.2. Embedding Layer
Similar to existing models, our approach utilizes an embedding layer to map item IDs to a high-dimensional space. This embedding layer uses a learnable embedding matrix $\bm{E}\in\mathbb{R}^{|\mathcal{V}|\times D}$, where $D$ is the embedding dimension. By applying the embedding layer to the input item sequence $\mathcal{S}_{u}$, we obtain the initial item embeddings $\mathcal{S}_{u}\bm{E}$. To enhance robustness and prevent overfitting, we incorporate both embedding dropout and layer normalization after retrieving the embeddings:
(3) $H$ $=\operatorname{LayerNorm}\left(\operatorname{Dropout}\left(\mathcal{S}_{u}\bm{E}\right)\right)\in\mathbb{R}^{n_{u}\times D}.$
During the training and inference stage, we merge multiple samples into one mini-batch, resulting in the input data $x\in\mathbb{R}^{B\times L\times D}$ for the subsequent Mamba block, where $B$ is the batch size and $L$ is padded sequence length.
@@ 3.3. Mamba Layer
Table 1. Recommendation performance. The best results are bold, and the second-best are underlined. The symbol * indicates a statistically significant improvement of Mamba4Rec over the best baseline with $p$-value ¡ 0.01.
@@
Mamba Block. As detailed in Algorithm 1, the Mamba block operates on an input $H_{i}\in\mathbb{R}^{B\times L\times D}$ with dimensions of batch size $B$, sequence length $L$, and hidden dimension $D$. It begins by performing linear projections on the input $H_{i}$ with an expanded hidden dimension of $D$ to obtain $H_{x}$ and $H_{z}$. These projections are then processed through a 1D convolution and a SiLU (elfwing2018sigmoid) activation. The core of the block involves a selective state space model with parameters discretized based on the input. This discretized SSM, along with $H^{\prime}_{x}$, generates the state representation $H_{y}$. Finally, $H_{y}$ is combined with a residual connection from $H_{z}$ after applying SiLU, and a final linear projection delivers the final output $H_{o}$. Overall, the Mamba block leverages input-dependent adaptations and selective SSM to process sequential information effectively. The parameters of the Mamba block consist of an SSM state expansion factor $N$, a kernel size $K$ for the convolution, and a block expansion factor $E$ for input and output linear projections. Selecting the expansion factors $N$ and $E$ involves a tradeoff between (capturing complex relationships) and training efficiency (higher $N,E$ increase computational cost).
The Mamba block addresses the prevalent effectiveness-efficiency dilemma existing in RNNs and Transformers. The recurrent SSM in the Mamba block operates in linear time with respect to the sequence length, making it as efficient as RNNs for inference, achieved by discretized parameters. Additionally, the selection mechanism makes the parameters of SSM input-dependent. This enables Mamba to selectively remember or forget information, crucial for capturing complex dependencies within sequential data. With the structured state matrix, it outperforms self-attention blocks in modeling long user behavioral sequences.
@@
1:Input: $H_{i}:(B,L,D)$
2:Output: $H_{o}:(B,L,D)$
3:$H_{x},H_{z}:(B,L,ED)\leftarrow\operatorname{Linear}(H_{i})$triangleright Linear projections of input
4:$H^{\prime}_{x}:(B,L,ED)\leftarrow\operatorname{SiLU}(\operatorname{Conv1d}(H_{x}))$
@@
7:$\Delta:(B,L,D)\leftarrow\operatorname{Softplus}(\text{Parameter}+\operatorname{Broadcast}(\operatorname{Linear}(H^{\prime}_{x})))$
@@
9:$H_{y}:(B,L,ED)\leftarrow\text{SelectiveSSM}(\overline{\bm{A}},\overline{\bm{B}},\bm{C})(H^{\prime}_{x})$
10:$H_{y^{\prime}}:(B,L,ED)\leftarrow H_{y}\otimes\operatorname{SiLU}(H_{z})$
11:$H_{o}:(B,L,D)\leftarrow\operatorname{Linear}(H_{y^{\prime}})$triangleright Final linear projection
12:return $H_{o}$
@@ 3.4. Prediction Layer
(5) $\hat{y}=\operatorname{Softmax}\left(h\bm{E}^{\top}\right)\in\mathbb{R}^{|\mathcal{V}|},$
@@ 4.1. Experimental Setup
Video-Games $24,303$ $10,673$ $231,780$ $9.5$
Baselines. We compare Mamba4Rec against several baseline methods, which include BPR-MF (rendle2012bpr), CNN-based Caser (tang2018personalized), RNN-based models like GRU4Rec (hidasi2015session) and NARM (li2017neural), Transformer-based models such as SASRec (kang2018self) and BERT4Rec (sun2019bert4rec), and LRURec (yue2024linear) based on linear recurrent units (orvieto2023resurrecting).
@@
Implementation Details. In the default architecture of our model, a single Mamba layer is employed without position embeddings. The Adam optimizer (kingma2014adam) is used with a learning rate of 0.001. The training batch size is 2048, and the evaluation batch size is 4096. All models use an embedding dimension of 64. To address the sparsity of the Amazon datasets, a dropout rate of 0.4 is used, compared to 0.2 for MovieLens-1M. The maximum sequence length is set proportionally to the mean number of actions per user: 200 for MovieLens-1M and 50 for Amazon datasets. We follow RecBole (zhao2021recbole) for further implementation details. For the parameters of the Mamba block, the SSM state expansion factor is 32, the kernel size for 1D convolution is 4, and the block expansion factor for linear projections is 2.
@@ 4.2. Overall Performance
Table 1 demonstrates the overall recommendation performance. Among sequential recommendation baseline models, Transformer-based models (i.e., SASRec, BERT4Rec) and LRU-based models (i.e., LRURec) generally outperform RNN-based models. Notably, by incorporating the selective SSMs, Mamba4Rec generally achieves superior performance compared to all the baselines across sparse and dense datasets, as well as varying maximum sequence lengths, revealing its strength in accurately capturing user preferences and recommending the proper items. The results highlight the potential of state space models as a compelling alternative to other existing user behavior modeling operators (e.g., RNNs, Transformers, and LRUs) for sequential recommendation.
@@ 4.3. Model Efficiency
Table 3 details the efficiency performance of Transformer-based models, LRURec and Mamba4Rec, all measured on a single Nvidia A5000 GPU with 24GB memory. SASRec, BERT4Rec, and LRURec are configured with their default setting of two layers. Considering the tradeoff between efficiency and effectiveness, we use the model with a single Mamba layer instead of models with only a Mamba block or stacked Mamba layers on the long-sequence MovieLens-1M dataset. Mamba4Rec exhibits superior performance in terms of GPU memory usage, training time per epoch, and inference time per batch. Additionally, it demonstrates faster convergence, reaching its best validation performance around epoch 50, compared to LRURec (epoch 55), SASRec (epoch 70), and BERT4Rec (epoch 180). These results suggest that Mamba4Rec achieves superior efficiency and reduces GPU memory cost, particularly on long-sequence datasets when compared to Transformer-based models and LRURec.
@@
LRURec 16.69GB 319.08s 0.37s
Mamba4Rec 5.01GB 56.13s 0.11s
Table 4. Ablation analysis on ML-1M and Amazon-Beauty.